expand the following log[6x^2y/Z^3}
Answer by lawdej(58)

log[6x^2y/Z^3}=log(6x^2y) -log(z^3)
=log(6)+log(x^2)+log(y) -log(z^3)
=log(6)+2log(x)+log(y) -3log(z)